You have a credit card with a balance of $14,800 and an APR of 18.4 percent compounded monthly. You have been making monthly payments of $280 per month, but you have received a substantial raise and will increase your monthly payments to $355 per month. How many months quicker will you be able to pay off the account?
Answer #1

Calculating Time Period with $280 per month,

Using TVM Calculation,

N = [PV = 14,800, PMT = -280, FV = 0, I = 0.184/12]

N = 109.30 months

Calculating Time Period with $355 per month,

Using TVM Calculation,

N = [PV = 14,800, PMT = -355, FV = 0, I = 0.184/12]

N = 67 months

Reduction in Time Period = 109.30 - 67

Reduction in Time Period = 42.30 months